OpenStudy (anonymous):

What is the molality of acetic acid (C2H4O2) in an aqueous solution that is 15.00% acetic acid by mass?

Parth (parthkohli):

Let there be 100 g of this aqueous solution. Then there's 15 g of acetic acid and 85 g of water. In one mole of acetic acid, there's 12*2 + 1*4 + 16*2 = 60 g. So here, we've got 0.25 moles of solute in 0.085 kg of water. Divide those two to get the molality, which is 0.25/0.085 = 100/34.

OpenStudy (anonymous):

Thank you very much!